# SiDiaC: Sinhala Diachronic Corpus

### v1

@@ Abstract
SiDiaC, the first comprehensive Sinhala Diachronic Corpus, covers a historical span from the 5th to the 20th century CE. SiDiaC comprises 58k words across 46 literary works, annotated carefully based on the written date, after filtering based on availability, authorship, copyright compliance, and data attribution. Texts from the National Library of Sri Lanka were digitised using Google Document AI OCR, followed by post-processing to correct formatting and modernise the orthography. The construction of SiDiaC was informed by practices from other corpora, such as FarPaHC, particularly in syntactic annotation and text normalisation strategies, due to the shared characteristics of low-resourced language status. This corpus is categorised based on genres into two layers: primary and secondary. Primary categorisation is binary, classifying each book into Non-Fiction or Fiction, while the secondary categorisation is more specific, grouping texts under Religious, History, Poetry, Language, and Medical genres. Despite challenges including limited access to rare texts and reliance on secondary date sources, SiDiaC serves as a foundational resource for Sinhala NLP, significantly extending the resources available for Sinhala, enabling diachronic studies in lexical change, neologism tracking, historical syntax, and corpus-based lexicography.
@@ 1 Introduction
Sola lingua bona est lingua mortua [^1] Latin: The only good language is a dead language.; given that all languages that are in use, evolve with a gradual process of linguistic change over time, proposed to have originated from an initially gestural communication system Corballis (2017). Factors affecting this complex evolution include cultural influences, which drive irregular word meaning shifts through phenomena such as new technologies (e.g., cell to mean cell phone in addition to prison cell) or community-specific vernaculars (e.g., gay to mean homosexual in addition to carefree). These cultural shifts often impact nouns more significantly. Conversely, regular linguistic processes, such as subjectification (e.g., actually shifting from objective to subjective usage) or grammaticalisation (e.g., promise undergoing rich changes), cause more predictable semantic changes and tend to affect verbs, adjectives, and adverbs more readily Hamilton et al. (2016).
The Sinhala language is an Indo-European language, which possesses a rich and diverse literary heritage that has developed over the course of several millennia, with its origins tracing back to between the 3rd and 2nd centuries BCE  de Silva (2025). This language has undergone significant evolution and transformation throughout its history, resulting in the form of modern Sinhala that we engage with today. Sinhala is spoken as L1 by approximately 16 million people, primarily located on the island of Sri Lanka de Silva (2025). The Sinhala script, which is unique to the language, descends from the Indian Brahmi script Fernando (1949); De Mel et al. (2025). Sinhala is classified as a lower-resourced language (Category 02) according to the criteria presented by Ranathunga and de Silva (2022).
@@ 2.1 LatinlSE
McGillivray and Kilgarriff (2013) introduced LatinlSE, a 13-million-word historical Latin corpus developed for the Sketch Engine [^3] https://www.sketchengine.eu/, a leading corpus query tool. Covering an extensive 22-century period from the 2nd Century BCE to the 21st Century CE, LatinlSE is equipped with detailed metadata, including author, title, genre, era, date, and century.
The methodology for creating this corpus involved gathering texts from various online digital libraries, such as LacusCurtius [^4] https://penelope.uchicago.edu/Thayer/E/Roman/Texts/, IntraText [^5] https://www.intratext.com/, and Musisque Deoque [^6] https://www.mqdq.it/. This process ensured a broad classification of genres as prose and poetry, and the texts were converted into a verticalized format while preserving their metadata. A significant aspect of the creation process was the automatic linguistic annotation using advanced NLP tools. This included lemmatisation with the PROIEL [^7] https://www.hf.uio.no/ifikk/english/research/projects/proiel/ project’s morphological analyser, complemented by Quick Latin [^8] https://www.quicklatin.com/ for unrecognised forms. Part-of-Speech (POS) tagging was achieved by training TreeTagger Schmid (1999) on existing Latin treebanks, including the Index Thomisticus Treebank [^9] https://itreebank.marginalia.it/, the Latin Dependency Treebank Bamman and Crane (2006), and the PROIEL project’s Latin treebank Haug and Jøhndal (2008). This training helps disambiguate analyses and assign the most likely lemma and POS to each token in context. This comprehensive dataset allows users to perform sophisticated searches based on lemmas, POS, and context, facilitating the study of shifts in word meanings over time.
@@ 2.2 IcePaHC and FarPaHC
The Icelandic Parsed Historical Corpus (IcePaHCRögnvaldsson et al. (2012) is a one-million-word parsed historical corpus of Icelandic, spanning from the late 12th century to the early 21st century. But more relevant to our work in this study is The Faroese Parsed Historical Corpus (FarPaHC), a syntactically annotated corpus of Faroese historical texts, that is presented as a spin-off of IcePaHC. The reason for this relevance is that, according to Ranathunga and de Silva (2022), Faroese also belongs to Category 02, similar to Sinhala. The FarPaHC corpus has 53,000 words.
@@ 2.3 Other Historical Corpora
Pettersson and Borin (2019) provides a comprehensive survey on existing diachronic and historical corpora. The work by Keersmaekers and Van Hal (2024) presents a case study demonstrating how large-scale automated parsing of Greek papyri can create richly annotated diachronic resources. Chen and Liu (2025) have created a Chinese corpus from the last 30 years of news articles on land usage. Even the corpora in higher-resourced languages such as DIAKORP Kučera et al. (2015) (Czech), ARCHER Biber et al. (1994), and  Davies (2012) [^13] https://www.english-corpora.org/coha/ (English), and DTA Geyken et al. (2011) and GerManC Scheible et al. (2011) (German) differ in size, balance, annotation depth, and access models. DIAKORP offers seven centuries of Czech texts, though it lacks linguistic annotation. ARCHER samples English registers across four centuries in 50-year intervals, while COHA spans two centuries of American English with lemmatisation and POS tagging.
Penn Parsed Corpora of Historical English Taylor and Kroch (1994) (PPCHE) and SRCMF Stein and Prévost (2013) (Old French) are similar to FarPaHC in the sense that they, too, are manually annotated corpora which provide syntactic analyses suitable for structural studies. PPCHE, in particular, has influenced corpora in other languages through its Penn-Helsinki annotation scheme, facilitating cross-linguistic comparison. Similarly, the PROIEL treebank family Eckhoff et al. (2018) extends such comparisons to some Indo-European languages via aligned New Testament translations.
ReM Klein and Dipper (2016) (Middle High German), RIDGES Odebrecht et al. (2017) (German-Science), and the Swedish Culturomics Gigaword corpus Eide et al. (2016), offer layered annotation or harmonised spellings for OCR quality control. While all corpora use some metadata scheme to provide critical contextual information such as date, genre, region, and authorship, beyond that, the metadata coverage varies widely. However, it can be noted that TEI-based [^14] Text Encoding Initiative (TEI) guidelines metadata schemes are popular among European language corpora.
@@ 3 Methodology
In this section, we describe the methodology used to create this dataset from the ground up. The process involved careful attention to detail at every stage, from planning to the final presentation, ensuring that the data is valid and of high quality. The procedure included addressing copyright laws in Sri Lanka, acquiring data, extracting text, and performing post-processing and formatting of the data as shown in Figure 1.
@@ 3.2 Annotation by Written Date
The issue date of the identified books was clearly stated in the digital repository at Natlib. However, this does not imply that the books were actually written during those specified dates. In fact, a book could have been written centuries earlier, while the printed version was released much later. Therefore, a comprehensive analysis was conducted to ensure that the written year of each book was accurately represented, ensuring that the resulting SiDiaC dataset accurately represents a proper diachronic corpus.
Upon the recommendation of experts in Sinhala linguistics, we identified a comprehensive book on Sinhala literature that claims to encompass literature information from its inception until 1994 CE Sannasgala (2015). This text served as the primary reference for the establishment of the respective date anchors, employing both time periods and specific years as outlined. The date ranges identified in Sannasgala (2015) correspond either to the period during which the book was authored or to the time period in which the author lived.
@@ 3.4 Data Filtration
Our first limitation was the availability of scanned copies at the digital repository of Natlib, which restricted us to 221 books. Additionally, we were only able to determine the written dates or periods for 65 of the 233 books. Taking both the availability of scanned copies and the accessibility of written dates into account, the number of selected records was reduced to 59. We then further refined this selection following the copyright law based filtering process we discussed, resulting in a final total of 46 books as presented in Table 6.
@@ 3.5 Text Extraction using OCR
In the comparative study conducted by Jayatilleke and de Silva (2025), five different OCR engines were analysed for text extraction using a synthetically created image-text dataset for Sinhala. Based on this study, we identified two standout OCR engines: Google Document AI [^18] https://cloud.google.com/document-ai/ and Surya [^19] https://github.com/VikParuchuri/surya, with Surya being reported to outperform all other systems compared. However, during our text extraction process, we found that, under realistic conditions (unlike the synthetic conditions used in the study), Google Document AI provided more accurate results.
@@ 3.6 Post-Processing Extracted Text
While language understanding was not a strict requirement for addressing these formatting-related factors, all manual post-processing procedures were carried out by the authors using a human-in-loop strategy Lamba and Madhusudhan (2023). This approach involved correcting formatting errors within a single window that contained both page scans and editable transcripts Christy et al. (2017). The authors responsible for these corrections are native Sinhala speakers. The post-processing steps applied, along with examples, are further discussed and illustrated in Appendix A.
@@ 3.7 Creation of Metadata Files
The dataset consisted of folders, each dedicated to a specific book. Within each folder, there is a text file along with a metadata file. The metadata files contain information such as the title and author names in both Sinhala and romanised forms, as well as the genre, issue date, written date, and the OCR confidence level for each particular book. Most of these information fields were identified through the LatinlSE corpus McGillivray and Kilgarriff (2013). The overall composition of the SiDiaC corpus, including folder and file level examples, is illustrated in Figure 4.
@@
The title of each book was consistently provided. When known, the authors’ names were included; if the authors were unknown, they were labelled as "unknown." The issued date corresponds to the published year as listed in the digital repository of Natlib, while the written date was determined by referring to Sannasgala (2015), as explained in the section 3.2.
The genres of the books were selected based on the details provided by Sannasgala (2015), as well as the content evaluated by authors who are native Sinhala speakers. The classification process occurs at two levels. The primary level is broad and divides the books into two categories: ‘Fiction’ and ‘Non-Fiction.’ The secondary level is more specific, categorising the content of the books into five distinct classes: religious, history, poetry, language, and medical. This approach was inspired by the methodologies followed in IcePaHC and DIAKORP Rögnvaldsson et al. (2012); Kučera et al. (2015) corpora to ensure diverse genres. It is important to note that the first level of categorisation was applied to all documents, while the second level of categorisation was applied only to the books that fell into the selected specific categories.
@@ 4.1 OCR Performance
    Text Modernisation: The historical development of the Sinhala language has resulted in various eras of syntax being linguistically equivalent but different in grapheme representation Nandasara and Mikami (2016). It is clear that Document AI adjusts to generate text in modern Sinhala syntax, ensuring a consistent and unified syntax throughout the dataset. Given that the change is only at the grapheme level, this does not violate the syntactic or semantic properties of the word.
@@
    Morpheme Segmentation: In Historical Sinhala, certain words are combined without spaces, forming a closed compound Gaikwad and Saini (2024). This phenomenon was accurately identified by Document AI, which effectively performs morpheme segmentation.
@@ 4.2 Evaluation of Metadata
Assuming that the books with specified date ranges are attributed to the upper bound year, an analysis of the number of books per century was conducted, as illustrated in Table 1. The analysis reveals that the distribution of books in the corpus is heavily skewed toward the 20th century, with 28 out of 46 records originating after the 18th century. This trend may largely be attributed to the introduction of the printing press to Sri Lanka by the Dutch in 1737, which thereafter popularised book printing in the country Wickremasuriya (1978); Nandasara and Mikami (2016).
In the first level of genre classification between fiction and non-fiction, it is evident that there are more non-fiction books than fiction books in the corpus. At the second level of genre classification, religious texts and poetry dominate among the five categories. This predominance is largely due to the close relationship between Sinhala literary culture and Theravada Buddhism, which provided both subjects and a framework for preserving texts. Additionally, the influence of Sanskrit kavya traditions and courtly patronage, which valued literary artistry and prestige, also played a significant role Hallisey (2003).
@@
The OCR confidence levels are extremely high, with an average of 96.84% across all books and a minimum confidence score of 85.53%. Despite these encouraging figures, it is clear that Document AI encountered various types of errors, which we largely addressed during the post-processing phase as discussed in Appendix A. The accurate identification of characters and words likely contributes to these strong confidence scores; however, most errors appear to arise from the challenges presented by complex content formats.

### v2

@@ Abstract
SiDiaC, the first comprehensive Sinhala Diachronic Corpus, covers a historical span from the 5th to the 20th century CE. SiDiaC comprises 58k words across 46 literary works, annotated carefully based on the written date, after filtering based on availability, authorship, copyright compliance, and data attribution. Texts from the National Library of Sri Lanka were digitised using Google Document AI OCR engine, followed by post-processing to correct formatting and modernise the orthography. The construction of SiDiaC was informed by practices from other corpora, such as FarPaHC, particularly in syntactic annotation and text normalisation strategies, due to the shared characteristics of low-resourced language status. This corpus is categorised based on genres into two layers: primary and secondary. Primary categorisation is binary, classifying each book into Non-Fiction or Fiction, while the secondary categorisation is more specific, grouping texts under Religious, History, Poetry, Language, and Medical genres. Despite challenges including limited access to rare texts and reliance on secondary date sources, SiDiaC serves as a foundational resource for Sinhala NLP, significantly extending the resources available for Sinhala, enabling diachronic studies in lexical change, neologism tracking, historical syntax, and corpus-based lexicography.
@@ 1 Introduction
Sola lingua bona est lingua mortua [^1] Latin: The only good language is a dead language.; given that all languages that are in use, evolve with a gradual process of linguistic change over time, proposed to have originated from an initially gestural communication system corballis2017evolution. Factors affecting this complex evolution include cultural influences, which drive irregular word meaning shifts through phenomena such as new technologies (e.g., cell to mean cell phone in addition to prison cell) or community-specific vernaculars (e.g., gay to mean homosexual in addition to carefree). These cultural shifts often impact nouns more significantly. Conversely, regular linguistic processes, such as subjectification (e.g., actually shifting from objective to subjective usage) or grammaticalisation (e.g., promise undergoing rich changes), cause more predictable semantic changes and tend to affect verbs, adjectives, and adverbs more readily hamilton-etal-2016-cultural.
The Sinhala language is an Indo-European language, which possesses a rich and diverse literary heritage that has developed over the course of several millennia, with its origins tracing back to between the 3rd and 2nd centuries BCE  de2025survey. This language has undergone significant evolution and transformation throughout its history, resulting in the form of modern Sinhala that we engage with today. Sinhala is spoken as L1 by approximately 16 million people, primarily located on the island of Sri Lanka de2025survey. The Sinhala script, which is unique to the language, descends from the Indian Brahmi script fernando1949palaeographical; de-mel-etal-2025-sinhala. Sinhala is classified as a lower-resourced language (Category 02) according to the criteria presented by ranathunga-de-silva-2022-languages.
@@ 2.1 LatinlSE
mcgillivray2013tools introduced LatinlSE, a 13-million-word historical Latin corpus developed for the Sketch Engine [^3] https://www.sketchengine.eu/, a leading corpus query tool. Covering an extensive 22-century period from the 2nd Century BCE to the 21st Century CE, LatinlSE is equipped with detailed metadata, including author, title, genre, era, date, and century.
The methodology for creating this corpus involved gathering texts from various online digital libraries, such as LacusCurtius [^4] https://penelope.uchicago.edu/Thayer/E/Roman/Texts/, IntraText [^5] https://www.intratext.com/, and Musisque Deoque [^6] https://www.mqdq.it/. This process ensured a broad classification of genres as prose and poetry, and the texts were converted into a verticalised format while preserving their metadata. A significant aspect of the creation process was the automatic linguistic annotation using advanced NLP tools. This included lemmatisation with the PROIEL [^7] https://www.hf.uio.no/ifikk/english/research/projects/proiel/ project’s morphological analyser, complemented by Quick Latin [^8] https://www.quicklatin.com/ for unrecognised forms. Part-of-Speech (POS) tagging was achieved by training TreeTagger schmid1999improvements on existing Latin treebanks, including the Index Thomisticus Treebank [^9] https://itreebank.marginalia.it/, the Latin Dependency Treebank bamman2006design, and the PROIEL project’s Latin treebank haug2008creating. This training helps disambiguate analyses and assign the most likely lemma and POS to each token in context. This comprehensive dataset allows users to perform sophisticated searches based on lemmas, POS, and context, facilitating the study of shifts in word meanings over time.
@@ 2.2 IcePaHC and FarPaHC
The Icelandic Parsed Historical Corpus (IcePaHCrognvaldsson-etal-2012-icelandic is a one-million-word parsed historical corpus of Icelandic, spanning from the late 12th century to the early 21st century. But more relevant to our work in this study is the Faroese Parsed Historical Corpus (FarPaHC), a syntactically annotated corpus of Faroese historical texts, that is presented as a spin-off of IcePaHC. The reason for this relevance is that, according to ranathunga-de-silva-2022-languages, Faroese also belongs to Category 02, similar to Sinhala. The FarPaHC corpus has 53,000 words.
@@ 2.3 Other Historical Corpora
pettersson2019characteristics provides a comprehensive survey on existing diachronic and historical corpora. The work by keersmaekers2024creating presents a case study demonstrating how large-scale automated parsing of Greek papyri can create richly annotated diachronic resources. chen2025administrative have created a Chinese corpus from the last 30 years of news articles on land usage. Even the corpora in higher-resourced languages such as DIAKORP kuvcera2015diakorp (Czech), ARCHER biber1994archer, COHA davies2012expanding [^13] https://www.english-corpora.org/coha/ (English), DTA geyken2011deutsche and GerManC scheible-etal-2011-gold (German) differ in size, balance, annotation depth, and access models. DIAKORP offers seven centuries of Czech texts, though it lacks linguistic annotation. ARCHER samples English registers across four centuries in 50-year intervals, while COHA spans two centuries of American English with lemmatisation and POS tagging. Penn Parsed Corpora of Historical English taylor1994penn (PPCHE) and SRCMF prevost2013syntactic (Old French) are similar to FarPaHC in the sense that they, too, are manually annotated corpora which provide syntactic analyses suitable for structural studies. PPCHE, in particular, has influenced corpora in other languages through its Penn-Helsinki annotation scheme, facilitating cross-linguistic comparison. Similarly, the PROIEL treebank family eckhoff2018proiel extends such comparisons to some Indo-European languages via aligned New Testament translations.
ReM klein2016handbuch (Middle High German), RIDGES odebrecht2017ridges (German-Science), and the Swedish Culturomics Gigaword corpus eide2016swedish, offer layered annotation or harmonised spellings for OCR quality control. While all corpora use some metadata scheme to provide critical contextual information such as date, genre, region, and authorship, beyond that, the metadata coverage varies widely. However, it can be noted that TEI-based [^14] Text Encoding Initiative (TEI) guidelines metadata schemes are popular among European language corpora.
@@ 3 Methodology
In this section, we describe the methodology used to create this dataset from the ground up. The process involved careful attention to detail at every stage, from planning to the final presentation, ensuring that the data were valid and of high quality. The procedure included addressing copyright laws in Sri Lanka, acquiring data, extracting text, and performing post-processing and formatting of the data as shown in Figure 1.
@@ 3.2 Annotation by Written Date
The issue date of the identified books was clearly stated in the digital repository at Natlib. However, this does not imply that the books were actually written during those specified dates. In fact, a book could have been written centuries earlier, while the printed version was released much later.
Document dating has become extensively recognised in computational sociology and studies within digital humanities ren-etal-2023-time; baledent-etal-2020-dating; hellwig-2020-dating. When compared to other dating tasks, historical text dating is more complex due to the absence of explicit temporal indicators (such as time expressions) that aid in determining the date a document was written toner2019language; baledent-etal-2020-dating; hellwig-2020-dating. It is clear that text dating, or the process of annotating the written date of a document, is an important task in diachronic studies ansari2023diachronic; ren-etal-2023-time; favaro-etal-2022-towards.
Therefore, a comprehensive analysis was conducted to ensure that the written year of each book was accurately represented, ensuring that the resulting SiDiaC dataset accurately represents a proper diachronic corpus.
Upon the recommendation of experts in Sinhala linguistics, we identified a comprehensive book on Sinhala literature that claims to encompass literature information from its inception until 1994 CE Sannasgala_2009. This text served as the primary reference for the establishment of the respective date anchors, employing both time periods and specific years as outlined. The date ranges identified in Sannasgala_2009 correspond either to the period during which the book was authored or to the time period in which the author lived.
@@ 3.4 Data Filtration
Our first limitation was the availability of scanned copies at the digital repository of Natlib, which restricted us to 221 books. Additionally, we were only able to determine the written dates or periods for 65 of the 233 books. Taking both the availability of scanned copies and the accessibility of written dates into account, the number of selected records was reduced to 59. We then further refined this selection following the copyright law based filtering process we discussed, resulting in a final total of 46 books as presented in Table 3 in Appendix A.
@@ 3.5 Text Extraction using OCR
In the comparative study conducted by jayatilleke2025zero, five different OCR engines were analysed for text extraction using a synthetically created image-text dataset for Sinhala. Based on this study, we identified two standout OCR engines: Google Document AI [^18] https://cloud.google.com/document-ai/ and Surya [^19] https://github.com/VikParuchuri/surya, with Surya being reported to outperform all other systems compared. However, during our text extraction process, we found that, under realistic conditions (unlike the synthetic conditions used in the study), Google Document AI provided more accurate results as shown in Appendix B.
@@ 3.6 Post-Processing Extracted Text
While language understanding was not a strict requirement for addressing these formatting-related factors, all manual post-processing procedures were carried out by the authors using a human-in-loop strategy lamba2023exploring. This approach involved correcting formatting errors within a single window that contained both page scans and editable transcripts christy2017mass. The authors responsible for these corrections are native Sinhala speakers. The post-processing steps applied, along with examples, are further discussed and illustrated in Appendix C.
@@ 3.7 Creation of Metadata Files
The dataset consisted of folders, each dedicated to a specific book. Within each folder, there is a text file along with a metadata file. The metadata files contain information such as the title and author names in both Sinhala and romanised forms, as well as the genre, issue date, written date, and the OCR confidence level for each particular book. Most of these information fields were identified through the LatinlSE corpus mcgillivray2013tools. Following the conventions of davies2012expanding and rognvaldsson-etal-2012-icelandic, we maintain a consistent metadata annotation method throughout the corpus without changing it across the centuries. The overall composition of the SiDiaC corpus, including folder and file level examples, is illustrated in Figure 4.
@@
The title of each book was consistently provided. When known, the authors’ names were included; if the authors were unknown, they were labelled as unknown. The issued date corresponds to the published year as listed in the digital repository of Natlib, while the written date was determined by referring to Sannasgala_2009, as explained in the section 3.2.
The genres of the books were selected based on the details provided by Sannasgala_2009, as well as the content evaluated by authors who are native Sinhala speakers. The classification process occurs at two levels. The primary level is broad and divides the books into two categories: ‘Fiction’ and ‘Non-Fiction.’ The secondary level is more specific, categorising the content of the books into five distinct classes: religious, history, poetry, language, and medical. This approach was inspired by the methodologies followed in IcePaHC and DIAKORP rognvaldsson-etal-2012-icelandic; kuvcera2015diakorp corpora to ensure diverse genres. It is important to note that the first level of categorisation was applied to all documents, while the second level of categorisation was applied only to the books that fell into the selected specific categories.
@@ 4.1 OCR Performance
    Text Modernisation: The historical development of the Sinhala language has resulted in various eras of syntax being linguistically equivalent but different in grapheme representation nandasara2016bridging. It is clear that Document AI adjusts to generate text in modern Sinhala syntax, ensuring a consistent and unified syntax throughout the dataset. Given that the change is only at the grapheme level, this does not violate the syntactic or semantic properties of the word.
@@
    Morpheme Segmentation: In Historical Sinhala, certain words are combined without spaces, forming a closed compound gaikwad2024Identification. This phenomenon was accurately identified by Document AI, which effectively performs morpheme segmentation.
@@ 4.2 Evaluation of Metadata
Assuming that the books with specified date ranges are attributed to the upper bound year, an analysis of the number of books per century was conducted, as illustrated in Table 1. The analysis reveals that the distribution of books in the corpus is heavily skewed toward the 20th century, with 28 out of 46 records originating after the 18th century. This trend may largely be attributed to the introduction of the printing press to Sri Lanka by the Dutch in 1737, which thereafter popularised book printing in the country wickremasuriya1978beginnings; nandasara2016bridging.
In the first level of genre classification between fiction and non-fiction, it is evident that there are more non-fiction books than fiction books in the corpus. At the second level of genre classification, religious texts and poetry dominate among the five categories. This predominance is largely due to the close relationship between Sinhala literary culture and Theravada Buddhism, which provided both subjects and a framework for preserving texts. Additionally, the influence of Sanskrit kavya traditions and courtly patronage, which valued literary artistry and prestige, also played a significant role hallisey2003works.
@@
The OCR confidence levels are extremely high, with an average of 96.84% across all books and a minimum confidence score of 85.53%. Despite these encouraging figures, it is clear that Document AI encountered various types of errors, which we largely addressed during the post-processing phase as discussed in Appendix C. The accurate identification of characters and words likely contributes to these strong confidence scores; however, most errors appear to arise from the challenges presented by complex content formats.